# Tracailer: An Efficient Trajectory Planner for Tractor-Trailer Vehicles in Unstructured Environments

### v1

@@ III-B Optimization Problem
[…]
@@
where $\boldsymbol{c}_{p}=[[\boldsymbol{c}_{x_{1}},\boldsymbol{c}_{y_{1}}]^{\text{T}},[\boldsymbol{c}_{x_{2}},\boldsymbol{c}_{y_{2}}]^{\text{T}},...,[\boldsymbol{c}_{x_{\text{M}}},\boldsymbol{c}_{y_{\text{M}}}]^{\text{T}}]^{\text{T}}\in\mathbb{R}^{6\text{M}\times 2}$,$\boldsymbol{c}_{\mathfrak{s}}=[\boldsymbol{c}_{\mathfrak{s}_{1}}^{\text{T}},\boldsymbol{c}_{\mathfrak{s}_{2}}^{\text{T}},...,\boldsymbol{c}_{\mathfrak{s}_{\text{M}}}^{\text{T}}]^{\text{T}}\in\mathbb{R}^{6\text{M}\times 1}$,$\boldsymbol{c}_{\theta}=[\boldsymbol{c}_{\theta_{1}}^{\text{T}},\boldsymbol{c}_{\theta_{2}}^{\text{T}},...,\boldsymbol{c}_{\theta_{\Omega}}^{\text{T}}]^{\text{T}}\in\mathbb{R}^{6\Omega\times\text{N}}$, $\boldsymbol{c}_{\theta_{k}}=[\boldsymbol{c}_{\theta_{11}},\boldsymbol{c}_{\theta_{12}},...,\boldsymbol{c}_{\theta_{\text{1N}}}]\in\mathbb{R}^{6\times\text{N}}$ are coefficient matrices. $\boldsymbol{C}=\{\boldsymbol{c}_{p},\boldsymbol{c}_{\mathfrak{s}},\boldsymbol{c}_{\theta}\}$. For the other optimization variables, $\boldsymbol{e}=[x_{0}^{\text{fina}},y_{0}^{\text{fina}},\theta_{0}^{\text{fina}},\boldsymbol{\theta}^{\text{fina}}]^{\text{T}}$, $T_{f}=\text{M}T_{p}=\Omega T_{\theta}$ denote the end state of the robot and total duration of the trajectory, respectively, where $\boldsymbol{\theta}^{\text{fina}}=[\theta_{1}^{\text{fina}},\theta_{2}^{\text{fina}},...,\theta_{\text{N}}^{\text{fina}}]$.
@@
Conditions (38$\sim$(39) denote the positive duration and the constraints brought by the introduction of $\mathfrak{s}(t)$ mentioned in the last subsection. Conditions (40$\sim$(42) are dynamic feasibility constraints, including avoiding jackknife[3], limitation of longitudinal velocity $v_{0}(t)$, longitude acceleration $a(t)$, latitude acceleration $a_{\text{lat}}(t)=v_{0}^{2}(t)\kappa(t)$, and curvature $\kappa(t)$, where $\delta\theta_{\text{max}},v_{\text{mlon}},a_{\text{mlon}},a_{\text{mlat}},\kappa_{\text{max}}$ are constants.
@@
$\boldsymbol{R}_{i}(t)$ $=\left[\begin{matrix}\cos\theta_{i}(t)&-\sin\theta_{i}(t)\\ \sin\theta_{i}(t)&\cos\theta_{i}(t)\end{matrix}\right]\quad\quad\quad i=1,...,\text{N}.$ (59)
@@ III-C Problem Simplification and Solving
$\vartheta_{di}^{\text{fina}}$ $=\text{L}_{c2}(\frac{\delta\theta_{\text{max}}+\theta_{di}^{\text{fina}}}{\delta\theta_{\text{max}}-\theta_{di}^{\text{fina}}})$ $i=1,2,...,N.$ (62)
@@
Then, we use Powell-Hestenes-Rockafellar Augmented Lagrangian method[42] (PHR-ALM) to solve the simplified problem, which is an effective method for solving large-scale problems and those with complex constraints. This method combines the principles of the Lagrangian method with an augmentation term to improve convergence properties, addressing both equality and inequality constraints by incorporating Lagrange multipliers and a penalty term that penalizes constraint violations. In this work, the unconstrained optimizer L-BFGS[43] is chosen as the inner loop iteration algorithm for PHR-ALM. It is a quasi-Newton method for solving large-scale unconstrained optimization problems with the advantage of using limited memory, which is particularly suitable for problems with a large number of variables, where storing the full Hessian would would be computationally prohibitive.
III-D Initial Value Acquisition
There are many methods that can be used to obtain initial values for trajectory optimization. Work[34] uses the A* algorithm search directly for all vehicles of the robot to obtain the path quickly, but since this path is kinematically infeasible, the optimization will take longer to converge to a good solution. Work[2] proposed an extended Hybrid-A* algorithm[12] that searches the trajectory in the space ×$\mathbb{R}\times SO(2)^{\text{N+1}}$ including the position of the tractor and yaw angles of all vehicles, which can obtain a better initial value. However, due to the large search space, it usually takes much more time, which leads to the whole system not meeting the requirement of real-time replanning in unknown environments for avoiding just-appearing obstacles. To balance the quality and efficiency, we propose a multi-terminal path search method based on Hybrid-A*[12] SE(2)-MHA, which searches only in $SE(2)$ space for the tractor and uses an approximate integration method to obtain the states of the trailers in searched paths corresponding to each terminal. Finally, we evaluate the quality of the paths using the weighted cost of the path length and the distance of each vehicle from the target region, choosing the one with the smallest value as the initial value for trajectory optimization. The pseudo code of the proposed method is given in Algorithm 1.
@@ IV-A Implementation details
Figure 3: Software system of the robot.
Fig. 3 illustrates the software system of the robot. In the perception module, we use the combination of LiDARs and reflective stickers to acquire the state of the robot, which includes the pose of the tractor and the yaw angle of each trailer. Specifically, the reflective stickers of the latter vehicle provides a high-intensity point cloud for the former LiDAR-installed vehicle. Then the relative pose of the two vehicles can be obtained by filtering this part of the point cloud and computing the normal vector of the fitted plane. Finally, using the pose in world frame output by the LiDAR-inertia Odometry (LIO) run on the LiDAR-installed vehicle, we can obtain the state of the robot by aligning it with other pose of vehicles. In this work, we adopt FAST-LIO2[45] as the LIO algorithm.
IV-B Real-World Experiments
@@
TABLE II: Statistics in a parking lot
@@ IV-C Cases Study
(b) Speed and curvature curves of optimized trajectories.
@@
In simulations, we first perform a qualitative and simple quantitative comparison for proposed method and OCP-STC[6] in a parking lot scenario with two different initial value acquisition methods, as shown in Fig. 7, where OCP-STC[6] is an efficient trajectory optimization method for tractor-trailer robots in unstructured environments, which uses Ipopt[51] as the solver and imposes obstacle avoidance constraints using safe travel corridors. The results using modified Hybrid-A* proposed in the work[2] as the initial value are shown in Fig. 7(a). Since the presence of narrow spaces and large turns in this scenario, it is difficult to provide good seeds from the front-end path for corridor generation for every vehicle in the robot, the OCP-STC[6] causes the robot to take bigger turns and does not make good use of the collision-free area at the edge of the parking space.
The results of the quantitative comparison are demonstrated in Table II. Since the SE(2)-MHA expands the states only in $SE(2)$ space for the tractor and does not perform collision check on the trailers, the time consumed by path search is reduced by almost an order of magnitude compared to Hybrid-A*. Although such an initial value leads to a longer optimization time, it reduces the total planning time. The corridor-based approach OCP-STC[6] requires completely collision-free initial values for corridor generation. Thus SE(2)-MHA cannot be used as an initial value. In addition, the corridors narrow the solution space of the problem, resulting in it not containing a better solution in this case, giving a longer trajectory length.
The second case is an ablation experiment to demonstrate the importance of slackened arc length in a simulated village, as shown in Fig. 8, where $v_{s}$ is the minimum speed constrained to avoid singularities. The black dashed lines in Fig. 8(b) indicate the boundaries of the constraints. As can be seen in Fig. 8(b), the smaller the $v_{s}$, the more significant the effect of singularity on trajectory optimization. Although the smaller the $v_{s}$, the more the optimizer is able to constrain the maximum curvature of the robot, it will make the speed of the tractor mutation at the initial and end states more severe, which is not in accordance with the real-world situation, and may cause difficulties for accurate trajectory tracking.
@@ IV-D Benchmark Comparisons
Figure 11: Benchmark Environments. The purple parts indicate obstacles. The height of the point cloud from low to high is indicated by the color from white to purple.
We carry out benchmark comparisons in a simulation environment of size $40m\times 40m$ with random triangular, quadrilateral, and pentagonal obstacles as shown in Fig. 11. We randomly initialize the initial state and target region of the robot in this environment. Among them, the initial state of the robot includes the $SE(2)$ state of the tractor and the yaw angle of each trailer. We use a rectangle as the target region, given that parking areas in the real world are usually of this shape. We set the length and width of the tractor trailer be $0.6m$ and $0.4m$, respectively. The length and width of the trailer are both set to be $0.4m$, the maximum speed is $2.0m/s$, the maximum longitudinal acceleration is $2.0m/s^{2}$, the minimum angle at which jacknife occurs is $1.47$rad. The wheelbase length of the tractor is set to $0.5m$, and the maximum steering angle of it is $0.7$ rad.
Two initial value acquisition methods and three optimization methods are tested in the above-mentioned environments. The initial value acquisition methods are the proposed SE(2)-MHA and Hybrid-A*, where the latter is the path search method proposed in the work[2], which does exploration over the entire ×$\mathbb{R}\times SO(2)^{\text{N+1}}$ state space. The optimization methods include OCP-STC[6], Proposed w.o. $\mathfrak{s}$ and Proposed w. $\mathfrak{s}$, where w.o. $\mathfrak{s}$ or w. $\mathfrak{s}$ is with or without using the slackened arc length as optimization variables. We compare these method in different environments with different density of obstacles when number of trailers N is 1, 2, or 3. Besides, tasks with different distance from the center of the rear wheels of the tractor to the center point of the target area ($3m\sim 10m$, $10m\sim 20m$, $20m\sim 40m$) are conducted separately. More than one hundred comparison tests are performed in each case. The parameters of all algorithms are finely tuned for the best performance.
More results on efficiency and quality are summarized in Table III. Where the tuple $(x,y,z)$ under environment complexity denotes the number of triangular, quadrilateral, and pentagonal random obstacles respectively. $l_{\text{traj}}$ denotes the mean length of the tractor trajectory, $t_{\text{d}}$ denotes the mean time duration of the optimized trajectory, and $\kappa$ denotes the mean curvature of the tractor, reflecting the smoothness of the trajectory. We set the front-end path search of greater than 5 seconds to count as failure. From the table, it can be seen that the success rate and computation time of SE(2)-MHA do not change much as the number of traiders becomes more. Its success rates are all greater than 98% and time consumption is all below 140 ms. This is attributed to the fact that it only searches in the $SE(2)$ space while utilizing the numerical integration approximation to obtain the trailers’ trajectory. On the contrary, since Hybrid-A* has to do the search and collision detection in the complete state space, the success rate decreases with the increase of the number of trailers under the time limit of 5 seconds. Meanwhile, it far exceeds SE(2)-MHA in terms of average time consumption, which is at least seven times more than the latter.
In the table, Hybrid-A* + Proposed w. $\mathfrak{s}$ performs best in most cases. Hybrid-A* searches the space more thoroughly, the initial values obtained tend to be better than SE(2)-MHA. In most cases, Proposed w. $\mathfrak{s}$ initialized with SE(2)-MHA performs second best and has a smaller gap with Hybrid-A* + Proposed w. $\mathfrak{s}$. Although the trajectory length and trajectory duration of Proposed w.o. $\mathfrak{s}$ are slightly better than OCP-STC[6] in most cases, the average curvature of the tractor trajectories is the largest in many cases due to the inability to handle the singularity. Since the strategy of generating corridors for each vehicle may lead to a smaller feasible space containing less optimal solutions, OCP-STC[6] does not perform as well as the proposed methods in terms of trajectory length and trajectory duration.
V Conclusion
In the future, we will further improve the adaptability of our algorithms to cope with more dynamic environments. In addition, integrating machine learning and embodied intelligence technologies to achieve smarter autonomous navigation, such as allowing tractor-trailer robots to acquire semantic information using GRB cameras to have more interaction with real-world environments, is also a promising direction to explore. Through these efforts, we believe that the application prospects for tractor-trailer robots in various transportation and logistics tasks will become even more expansive.
TABLE III: Benchmark Comparison in Different Cases

### v2

@@ III-B Optimization Problem
$\min_{\boldsymbol{C},\boldsymbol{e},T_{f}}$ $f(\boldsymbol{C},\boldsymbol{e},T_{f})=\int_{0}^{\Sigma_{j=1}^{\text{M}}\mathfrak{s}_{j}(T_{p})}\boldsymbol{j}_{p}(\mathfrak{s})^{\text{T}}\boldsymbol{W}_{p}\boldsymbol{j}_{p}(\mathfrak{s})d\mathfrak{s}$ (32)
$\quad\quad\quad\quad\quad\ +\int_{0}^{T_{f}}\boldsymbol{j}_{\theta}(t)^{\text{T}}\boldsymbol{W}_{\theta}\boldsymbol{j}_{\theta}(t)dt+\rho_{t}T_{f}$ (33)
@@
where $\boldsymbol{c}_{p}=[[\boldsymbol{c}_{x_{1}},\boldsymbol{c}_{y_{1}}]^{\text{T}},[\boldsymbol{c}_{x_{2}},\boldsymbol{c}_{y_{2}}]^{\text{T}},...,[\boldsymbol{c}_{x_{\text{M}}},\boldsymbol{c}_{y_{\text{M}}}]^{\text{T}}]^{\text{T}}\in\mathbb{R}^{6\text{M}\times 2}$$\boldsymbol{c}_{\mathfrak{s}}=[\boldsymbol{c}_{\mathfrak{s}_{1}}^{\text{T}},\boldsymbol{c}_{\mathfrak{s}_{2}}^{\text{T}},...,\boldsymbol{c}_{\mathfrak{s}_{\text{M}}}^{\text{T}}]^{\text{T}}\in\mathbb{R}^{6\text{M}\times 1}$$\boldsymbol{c}_{\theta}=[\boldsymbol{c}_{\theta_{1}}^{\text{T}},\boldsymbol{c}_{\theta_{2}}^{\text{T}},...,\boldsymbol{c}_{\theta_{\Omega}}^{\text{T}}]^{\text{T}}\in\mathbb{R}^{6\Omega\times\text{N}}$, $\boldsymbol{c}_{\theta_{k}}=[\boldsymbol{c}_{\theta_{11}},\boldsymbol{c}_{\theta_{12}},...,\boldsymbol{c}_{\theta_{\text{1N}}}]\in\mathbb{R}^{6\times\text{N}}$ are coefficient matrices. $\boldsymbol{C}=\{\boldsymbol{c}_{p},\boldsymbol{c}_{\mathfrak{s}},\boldsymbol{c}_{\theta}\}$. For the other optimization variables, $\boldsymbol{e}=[x_{0}^{\text{fina}},y_{0}^{\text{fina}},\theta_{0}^{\text{fina}},\boldsymbol{\theta}^{\text{fina}}]^{\text{T}}$, $T_{f}=\text{M}T_{p}=\Omega T_{\theta}$ denote the end state of the robot and total duration of the trajectory, respectively, where $\boldsymbol{\theta}^{\text{fina}}=[\theta_{1}^{\text{fina}},\theta_{2}^{\text{fina}},...,\theta_{\text{N}}^{\text{fina}}]$.
@@
[…]
@@
$\boldsymbol{R}_{i}(t)$ $=\left[\begin{matrix}\cos\theta_{i}(t)&-\sin\theta_{i}(t)\\ \sin\theta_{i}(t)&\cos\theta_{i}(t)\end{matrix}\right],\quad\quad\quad i=1,...,\text{N}.$ (59)
@@ III-C Problem Simplification and Solving
$\vartheta_{di}^{\text{fina}}$ $=\text{L}_{c2}\left(\frac{\delta\theta_{\text{max}}+\theta_{di}^{\text{fina}}}{\delta\theta_{\text{max}}-\theta_{di}^{\text{fina}}}\right)$ $i=1,2,...,N.$ (62)
@@
Then, we use Powell-Hestenes-Rockafellar Augmented Lagrangian method[45] (PHR-ALM) to solve the simplified problem, which is an effective method for solving large-scale problems and those with complex constraints. To give readers a more accurate understanding of our work, we give a proof of convergence of this algorithm in the Appendix -A and provide a quantitative analysis to show the degree to which constraints are satisfied.
Inspired by other works on robot trajectory optimization problems that also need to deal with equation constraints[41, 46], we set the trajectory of trailers have more pieces (i.e. $\Omega>\text{M}$) to make it easier for the optimizer to converge and better fit the kinematic constraints (34). Based on engineering experience, we found that the solver can achieve good efficiency and success when $\Omega=\text{ceil}(2.0\times\text{M})$, where function $\text{ceil}(x)$ serves to take the smallest integer not less than $x$.
@@
Figure 3: Software system of the robot. The perception module includes the global localization of the robot, its own state estimation, and traversability assessment. The planning and control module includes the proposed planning pipeline and trajectory tracking controller.
III-D Initial Value Acquisition
There are many methods that can be used to obtain initial values for trajectory optimization. Work[36] uses the A* algorithm search directly for all vehicles of the robot to obtain the path quickly, but since this path is kinematically infeasible, the optimization will take longer to converge to a good solution. Work[3] proposed an extended Hybrid-A* algorithm[14] that searches the trajectory in the space $\mathbb{R}\times SO(2)^{\text{N+1}}$ including the position of the tractor and yaw angles of all vehicles, which can obtain a better initial value. However, due to the large search space, it usually takes much more time, which leads to the whole system not meeting the requirement of real-time replanning in unknown environments for avoiding just-appearing obstacles.
To balance the quality and efficiency, we propose a multi-terminal path search method based on Hybrid-A*[14] SE(2)-MHA, which searches only in $SE(2)$ space for the tractor and uses an approximate integration method similar to that in work [3] to obtain the path and terminal position of the trailers in searched paths corresponding to each terminal. Finally, we evaluate the quality of the paths using the weighted cost of the path length and the distance of each vehicle from the target region, choosing the one with the smallest value as the initial value for trajectory optimization. Specifically, assuming the length of the tractor path is $l_{\text{tra}}$ and the distance between the terminal position of $i$-th trailer and the target convex area is $d_{it},i=1,2,\text{N}$, the cost of the path is $\text{Cost}_{traj}=w_{\text{r2}}l_{\text{tra}}+w_{\text{trailer}}\sum_{i=1}^{\text{N}}d_{it}$, where $w_{\text{r2}}$ and $w_{\text{trailer}}$ are the weights. In our algorithm, $d_{it}$ will be set to zero if the terminal position of the $i$-th trailer is within the target convex area. The pseudo code of the proposed method is given in Algorithm 1.
@@ III-D Initial Value Acquisition
IV Results
IV-A Implementation details
Fig. 3 illustrates the software system of the robot. In the perception module, we use the combination of LiDARs and reflective stickers[48, 49] to acquire the state of the robot, which includes the pose of the tractor and the yaw angle of each trailer. Specifically, the reflective stickers of the latter vehicle provides a high-intensity point cloud for the former LiDAR-installed vehicle. Then the relative pose of the two vehicles can be obtained by filtering this part of the point cloud and computing the normal vector of the fitted plane. We run LiDAR-inertia Odometry (LIO) on two LiDARs separately. Using the initial point cloud map constructed by the LiDAR on the tractor as initial map of the LIO on the trailer, we can align the coordinate system of these two LIOs. In this work, we adopt FAST-LIO2[50] as the LIO algorithm.
IV-B Real-World Experiments
@@
TABLE II: Key Parameter Settings for Simulation in Parking Lot. M.P. represents the proposed method. M.O. represents OCP-STC[7]. The specific names of abbreviations m.s., mu.s., b.p., l.s., m.c.t are “mem size”, “mu strategy”, “bound push”, “linear solver”, “max cpu time”, respectively.
Type Para. Description Setting
M.P.
$L_{q}$
Intervals for taking state points from the
front-end path as optimization variables
5 m
$K$
The number of points sampled to impose
constraints in each polynomial trajectory
16
m.s.
The number of corrections to approximate
the inverse hessian matrix in L-BFGS
64
$\delta_{\text{inner}}$ Desired convergence tolerance for L-BFGS 1.0-4
past Distance for delta-based convergence test 3
$N_{\text{iter}}$ Maximum iteration number of L-BGFS 10000
$\epsilon_{\text{cond}}$ Desired convergence tolerance for ALM 1.0-1
$N_{\text{out}}$ Maximum iteration number of ALM 100
M.O. $L_{fe}$
Intervals for initializing finite
elements from the front-end path
0.3 m
mu.s. Update strategy for barrier parameter adaptive
b.p.
Desired minimum absolute distance from
the initial point to bound
1.0e-8
l.s. The core linear solver for Ipopt MA97
m.c.t
Limit on CPU seconds that Ipopt can
use to solve one problem
200.0 s
TABLE III: Statistics in a parking lot
@@ IV-C Cases Study
In simulations, we first perform a qualitative and simple quantitative comparison for proposed method and OCP-STC[7] in a parking lot scenario with two different initial value acquisition methods, as shown in Fig. 7, where OCP-STC[7] is an efficient trajectory optimization method for tractor-trailer robots in unstructured environments, which uses Ipopt[56] as the solver and imposes obstacle avoidance constraints using safe travel corridors. In this case, we set $L_{0}=2.7m,\text{N}=3,L_{1}=L_{2}=L_{3}=4.0m.$ The width of the tractor and trailers $L_{w}=2.0m$. The constants related to kinematic constraints are set to $v_{\text{mlon}}=10.0m/s,a_{\text{mlon}}=5.0m/s^{2},a_{\text{mlat}}=10m/s^{2},\kappa_{\text{max}}=0.95m^{-1},\delta\theta_{\text{max}}=1.47\text{rad}$, respectively. As in the original paper[7], we use AMPL[57] to construct the optimization problem for OCP-STC[7]. Moreover, the list of key parameter settings for optimization algorithms is given in Table II. The results using modified Hybrid-A* proposed in the work[3] as the initial value are shown in Fig. 7(a). Since the presence of narrow spaces and large turns in this scenario, it is difficult to provide good seeds from the front-end path for corridor generation for every vehicle in the robot, the OCP-STC[7] causes the robot to take bigger turns and does not make good use of the collision-free area at the edge of the parking space.
The results of the quantitative comparison are demonstrated in Table III. Since the SE(2)-MHA expands the states only in $SE(2)$ space for the tractor and does not perform collision check on the trailers, the time consumed by path search is reduced by almost an order of magnitude compared to Hybrid-A*. Although such an initial value leads to a longer optimization time, it reduces the total planning time. The corridor-based approach OCP-STC[7] requires completely collision-free initial values for corridor generation. Thus SE(2)-MHA cannot be used as an initial value. In addition, the corridors narrow the solution space of the problem, resulting in it not containing a better solution in this case, giving a longer trajectory length.
@@
(b) Speed, curvature, tracking error curves of optimized trajectories.
@@
The second case is an ablation experiment to demonstrate the importance of slackened arc length in a simulated village, as shown in Fig. 8, where $v_{s}$ is the minimum speed constrained to avoid singularities. In this case, we set $L_{0}=2.6m,\text{N}=3,L_{1}=L_{2}=L_{3}=4.0m.$ The width of the tractor and trailers $L_{w}=2.8m$. The constants related to kinematic constraints are set to $v_{\text{mlon}}=10.0m/s,a_{\text{mlon}}=5.0m/s^{2},a_{\text{mlat}}=2m/s^{2},\kappa_{\text{max}}=0.99m^{-1},\delta\theta_{\text{max}}=1.47\text{rad}$, respectively. Moreover, the key parameters of the optimization algorithm are set the same as those given in Table II for methods with and without the introduction of slackened arc length $\mathfrak{s}$. The black dashed lines in Fig. 8(b) indicate the boundaries of the constraints. As can be seen in Fig. 8(b), the smaller the $v_{s}$, the more significant the effect of singularity on trajectory optimization. Although the smaller the $v_{s}$, the more the optimizer is able to constrain the maximum curvature of the robot, it will make the speed of the tractor mutation at the initial and end states more severe, which is not in accordance with the real-world situation, and may cause difficulties for accurate trajectory tracking. We introduce a model predictive control (MPC)-based controller[58] to measure the tracking error in the simulation. In our paper, the tracking error of a tractor-trailer robot is defined as the average tracking error of all vehicles in the robot. The statistical results are shown in Table IV and the curves in Fig. 8(b). Optimized trajectories without the use of $\mathfrak{s}$ lead to larger tracking errors due to abrupt changes in velocity or constraints violations during the initial and end phases of the trajectories. Especially when $v_{s}=1.0$, extremely discontinuous velocities cause the controller to almost diverge at the beginning.
TABLE IV: Tracking error statistics for ablation experiment.
Method Mean T.E. Max T.E. STD.
(m) (m) (m)
Proposed w.o. $\mathfrak{s}$
$v_{s}=0.01$
0.227 0.434 0.0959
Proposed w.o. $\mathfrak{s}$
$v_{s}=0.1$
0.191 0.366 0.0868
Proposed w.o. $\mathfrak{s}$
$v_{s}=1.0$
0.443 1.06 0.259
Proposed w. $\mathfrak{s}$ 0.133 0.254 0.0595
Figure 9: Physical simulation based on Gazebo. In Rviz, the purple parts indicate obstacles. The height of the point cloud from low to high is indicated by the color from white to purple.
@@ IV-D Benchmark Comparisons
We carry out benchmark comparisons in a simulation environment of size $40m\times 40m$ with random triangular, quadrilateral, and pentagonal obstacles as shown in Fig. 9. We randomly initialize the initial state and target region of the robot in this environment. Among them, the initial state of the robot includes the $SE(2)$ state of the tractor and the yaw angle of each trailer. We use a rectangle as the target region, given that parking areas in the real world are usually of this shape. We set the length and width of the tractor trailer be $0.6m$ and $0.4m$, respectively. The length and width of the trailer are both set to be $0.4m$, the maximum speed is $2.0m/s$, the maximum longitudinal acceleration is $2.0m/s^{2}$, the minimum angle at which jacknife occurs is 1.471.47rad. The wheelbase length of the tractor is set to $0.5m$, and the maximum steering angle of it is $0.7$ rad. The parameters of all algorithms are finely tuned for the best performance. Moreover, the key parameters for optimization algorithms are consistent with those in Table II, except that considering the scene and robot size, we set $L_{q}=1.0$ m, $K=8$, and $L_{fe}=0.1m$. As in Sec.IV-C, we follow the original paper[7] and deploy OCP-STC using AMPL[57].
@@
Two initial value acquisition methods and three optimization methods are tested in the above-mentioned environments. The initial value acquisition methods are the proposed SE(2)-MHA and Hybrid-A*, where the latter is the path search method proposed in the work[3], which does exploration over the entire $\mathbb{R}\times SO(2)^{\text{N+1}}$ state space. The optimization methods include OCP-STC[7], Proposed w.o. $\mathfrak{s}$ and Proposed w. $\mathfrak{s}$, where w.o. $\mathfrak{s}$ or w. $\mathfrak{s}$ is with or without using the slackened arc length as optimization variables. We compare these method in different environments with different density of obstacles when number of trailers N is 1, 2, or 3. Besides, tasks with different distance from the center of the rear wheels of the tractor to the center point of the target area ($3m\sim 10m$, $10m\sim 20m$, $20m\sim 40m$) are conducted separately. More than one hundred comparison tests are performed in each case.
@@
TABLE V: Benchmark Comparison in Different Cases